A secretary commits 0,1,2,3,4, or 5 encoding errors on a given hour with probabilities 0.09, 0.30, 0.25, 0.15, 0.10, and 0.06 respectively. Find the probability that on the next hour she will commit more than 5 errors.


1
Expert's answer
2022-02-23T15:10:03-0500

Let X be number of errors commited by secretary, then

"P(X>5)=1-P(X=0)-P(X=1)-P(X=2)-P(X=3)-P(X=4)-P(X=5)=1-0.09-0.3-0.25-0.15-0.1-0.06=0.05"